Why is the variance of a distrubtion of means smaller than the original popluation variance?
It is a result of the Central Limit Theorem.
What are the next set of numbers after 12 34 35 38 42?
Given any number it is easy to find a rule based on a polynomial of order 5+k such that the first five numbers are as listed in the question and the next k are the given "next" numbers.
However, use the rule
Un = (-6n4 + 83n3 - 411n2 + 874n - 468)/6
Accordingly, the next three numbers are 22, -71, -310
What is the probability of rolling and even number on an ordinary dice?
your set of even number x = {2,4,6}, which tells you that 3 out the 6 sides are even. Your probability is 3/6 = 1/2
What can undermine the value of using standard costing?
Imposing standard costs without communicating in an honest, candid manner will undermine much of the perceived value of such costing.
What evidence might suggest the data is approximately normally distributed or not?
If the pearson's skew is between 0.0 and 0.49, then it is a normal or non-problematic distribution. If it is above .50, then it is not a normal distribution.
In appearance:
The data should be symmtrical; the mode and median should be very close to the mean; around 65 % of the observations should be withing one standard deviation (SD) of the mean; 95% of observations with 2 SD of the mean and almost all within 3 sd.
Better evidence can be provided by the chi-square or Kolmogorov-Smirnoff goodness-of-fir tests.
There may also be evidence from the nature of the data. For example, the means of samples will be closer to being normally distributed (Central Limit Theorem) than the individual observations.
How large should our sample be?
That really depends on the specific circumstances - including the level of confidence you want to achieve.
How do you graph a line using polar coordinates?
This only works for y=mx, not y=mx+c.
theta = tan(m)
eg y=x theta = tan(1)
.: theta = 45
(23,56)
Is 13 decimeters closest to 1 meter?
No. 11 decimetres = 1.1 metre is closer. And 1.01 m is closer still. 1.001 m is even closer and so on.
What is the probability of rolling two dice and getting the same number?
1/6 is the answer
Another Answer:-
It is: 1/6*1/6 = 1/36
Will running more trials affect results of a sampling distribution?
Yes. It should reduce the variance.
How do you get a list of all four digit numbers?
It is actually very simple, just take 10 to the square root of four.
10^4